### v1

# Simplifying Traffic Anomaly Detection with Video Foundation Models

Abstract
Recent methods for ego-centric Traffic Anomaly Detection (TAD) often rely on complex multi-stage or multi-representation fusion architectures, yet it remains unclear whether such complexity is necessary. Recent findings in visual perception suggest that foundation models, enabled by advanced pre-training, allow simple yet flexible architectures to outperform specialized designs. Therefore, in this work, we investigate an architecturally simple encoder-only approach using plain Video Vision Transformers (Video ViTs) and study how pre-training enables strong TAD performance. We find that: (i) strong pre-training enables simple encoder-only models to match or even surpass the performance of specialized state-of-the-art TAD methods, while also being significantly more efficient; (ii) although weakly- and fully-supervised pre-training are advantageous on standard benchmarks, we find them less effective for TAD. Instead, self-supervised Masked Video Modeling (MVM) provides the strongest signal; and (iii) Domain-Adaptive Pre-Training (DAPT) on unlabeled driving videos further improves downstream performance, without requiring anomalous examples. Our findings highlight the importance of pre-training and show that effective, efficient, and scalable TAD models can be built with minimal architectural complexity. We release our code, domain-adapted encoders, and fine-tuned models to support future work: https://github.com/tue-mps/simple-tad.
1 Introduction
Figure 1: Traffic Anomaly Detection (TAD) performance on DADA-2000 [13]. Simple encoder-only models with strong pre-training (blue) are faster and more accurate than recent multi-component architectures (orange and red), see Tabs. 2 and 3.
Traffic risk estimation is fundamental to safe driving, as failures to anticipate danger can lead to life-threatening consequences. Autonomous vehicles must therefore assess potential hazards in real time, even under uncertain, dynamic, and unfamiliar conditions. A common formulation for this problem is the ego-centric traffic anomaly detection (TAD) task [48, 13], which aims to identify abnormal or dangerous events in a video stream captured by a vehicle‐mounted camera. Analyzing the top-performing TAD methods [51, 24, 32, 25], we find that they rely on specialized, complex architectures illustrated in Fig. 2(b) and (c): two-stage approaches [51, 34, 10], which combine a vision encoder with a temporal component, and multi-representation fusion approaches [24, 32, 25], which fuse additional representations, often generated by separate deep neural networks or model-based algorithms. While these complex designs have improved performance, their impact on efficiency has not been evaluated, even though this is crucial for TAD, where rapid detection is needed to enable timely action and prevent accidents.
Figure 2: Types of model architectures for TAD: simple encoder-only model (a), two-stage design (b), and multi-representation fusion architecture (c).
Considering that TAD is, in essence, a binary classification task on video, we turn to recent methods for general video classification. On standard video classification benchmarks [19, 15, 8, 36], Video Foundation Models (ViFMs) achieve state-of-the-art performance, predominantly Video Vision Transformers (ViTs) [11, 1], which rely on large-scale self- and weakly-supervised pre-training to learn expressive and transferable spatiotemporal representations, rather than on architectural inductive biases. Indeed, recent work in related visual perception tasks shows that strong pre-training reduces the need for downstream task-specific components [21, 41, 20]. We hypothesize that the same applies to TAD, such that a simple ViT-based ViFM can be effectively applied to this task and match or even outperform complex architectures. Since TAD depends on motion understanding, we investigate whether pre-training strategies that capture spatio-temporal structure are particularly effective.
To test our hypotheses, we evaluate multiple ViT-based ViFMs, sharing the same plain video ViT architecture but different pre-training, on two TAD datasets, DoTA [48] and DADA-2000 [13]. We adopt an encoder-only design, with a single linear layer on top of the Video ViT model, as illustrated in Fig. 2(a). While prior work has attempted the encoder-only design [10], it was found to be inferior [24, 32, 34]. We revisit this design using stronger pre-training. Similar to prior work, we assess in-domain and out-of-domain generalization, and unlike prior work, we also assess computational efficiency of the models, and compare to the best-performing specialized TAD methods.
We confirm our hypothesis by showing that strong pre-training enables a plain video ViT, used as an encoder-only model for TAD, to match and even surpass state-of-the-art methods while also being significantly more efficient, as shown in Fig. 1 and Tab. 2. Interestingly, performance on standard video classification benchmarks does not correlate with TAD. Our comparison of pre-trained models shows that weak supervision from language and full supervision from class labels are effective on standard benchmarks but less so for TAD, likely because they promote appearance-focused features that generalize poorly to anomalous motion [45]. In contrast, self-supervised learning with Masked Video Modeling (MVM), which trains the model to reconstruct missing spatio-temporal regions using both spatial structure and temporal continuity, proves most effective for TAD. Models pre-trained with this objective achieve state-of-the-art performance at their respective model size, as shown in Tab. 3.
Next, motivated by the scarcity of labeled data for TAD and the abundance of unlabeled driving videos, we explore whether we can leverage self-supervised learning to better adapt an off-the-shelf ViFM to the downstream domain. Specifically, we apply Domain-Adaptive Pre-Training (DAPT) [16] using the Video Masked Autoencoding (MAE) [38] approach. We find that MAE-based DAPT, even when applied at relatively small scale compared to the preceding generic pre-training, significantly boosts the performance even further, particularly for smaller models, as shown in Fig. 3. Importantly, we find that including abnormal driving examples for DAPT is not necessary, as shown in Tab. 4, which is valuable given the difficulty of collecting them at scale.
We summarize our main contributions as follows:
  1. 1.
    We show that a plain encoder-only Video ViT, when equipped with strong pre-training, outperforms all prior specialized architectures for TAD, while also being significantly more efficient.
  2. 2.
    We compare pre-training strategies and find that self-supervised learning with a Masked Video Modeling (MVM) objective is most effective, outperforming both weakly- and fully-supervised alternatives.
  3. 3.
    We demonstrate that Domain-Adaptive Pre-training with MVM leads to measurable performance gains on TAD, even when applied at a small scale to domain-relevant but anomaly-free data.
Moreover, since every TAD pipeline relies on its visual encoder, our results offer clear guidance for selecting effective pre-training strategies, enabling future methods to detect traffic anomalies more accurately and robustly.
2 Related work
2.1 Traffic Anomaly Detection
Traffic Anomaly Detection (TAD) is typically framed as a binary video classification task, where the goal is to detect potentially dangerous or abnormal events in traffic scenarios from the egocentric viewpoint of a vehicle-mounted camera. While related to the broader field of Video Anomaly Detection (VAD), which commonly targets static-camera surveillance settings, TAD poses unique challenges due to ego-motion, frequent occlusions, and the dynamic interaction of agents in complex driving environments.
Before the availability of annotated datasets for TAD, earlier approaches often relied on unsupervised reconstruction or prediction of video frames to flag anomalies using temporal autoencoders [29, 43], future frame prediction with spatial, temporal, and adversarial losses [26], or spatiotemporal tubelet modeling with ensemble scoring [48]. Some works also explored the use of synthetic data for training [35, 22].
The introduction of large-scale driving anomaly datasets [48, 13] with comprehensive annotations has enabled more active development of fully-supervised methods and led to substantial improvements in detection performance. As shown in Fig. 2, we categorize TAD methods into three classes based on their architectural complexity, which we detail below.
Encoder-only design, (Fig. 2, a).  Since TAD is a binary classification task, a minimal solution consists of a feature encoder followed by a linear classifier, without additional task-specific modules. Prior work [10] shows that such designs can be effective in related tasks, where R(2+1)D [39] and ViViT [1] demonstrate considerably strong performance. However, several recent studies report underwhelming results for encoder-only ablation variants of their methods [24, 32]. Our evaluation of R(2+1)D and ViViT models on standard TAD benchmarks (see Tab. 2) reveals a substantial performance gap between these encoder-only models and current state-of-the-art methods.
Two-stage design, (Fig. 2, b).  Two-stage methods combine a visual encoder with a separate temporal module. VidNeXt [10] pairs a ConvNeXt [27] backbone with a non-stationary transformer (NST) to model both stable and dynamic temporal patterns, evaluating and introducing a new dataset CycleCrash [10] for the related task of collision prediction. Its ablations, ConvNeXt+VT and ResNet+NST, also yield strong results. MOVAD [34] uses a VideoSwin Transformer [28] as a short-term memory encoder over several frames, followed by an LSTM-based long-term module, achieving state-of-the-art performance for TAD.
Multi-representation fusion design, (Fig. 2, c).  Fusion-based models, which currently report state-of-the-art performance in TAD, explicitly combine multiple information sources. TTHF [24] augments the CLIP [33] framework with a high-frequency motion encoder and a cross-modal fusion module to align motion features with textual prompts. PromptTAD [32] extends MOVAD [34] by incorporating bounding box prompts via instance- and relation-level attention, enhancing object-centric anomaly localization. ISCRTAD [25] integrates agent features (e.g., appearance, trajectory, depth) using graph-based modeling and fuses them with scene context through contrastive multimodal alignment for robust anomaly detection.
In this work, we follow the simple encoder-only design and explore whether strong pre-training can compensate for the lack of task-specific architectural inductive biases. We demonstrate that, when equipped with rich priors from large-scale self-supervised pre-training, such models can achieve state-of-the-art performance, while remaining architecturally simple and highly efficient.
2.2 Video Foundation Models
Year Model Stage Type Objective Supervision
2021 ViViT [1] Stage 1 FSL Classification Class labels
2022 VideoMAE [38] Stage 1 SSL Masked Autoencoder Video frame pixels
2023 MVD [45] Stage 1 SSL Masked feature distillation High-level features of VideoMAE and ImageMAE teachers
2023 VideoMAE2 [44] Stage 1 SSL Dual MAE Video frame pixels
Stage 2 FSL Classification Class labels
Stage 3 FSL Logit distillation Logits of larger VideoMAE2 after stage 2
2024 InternVideo2 [46] Stage 1 WSL+SSL Unmasked feature distillation Features of VideoMAE2 and a vision-language encoder
Stage 2 WSL Feature distillation + Contrastive Features of audio and text encoders + video, text, audio
Stage 3 WSL+SSL Feature distillation Features of InternVideo2 after stage 2 across multiple depths
Table 1: Overview of ViFMs. For models trained via distillation, we denote the supervision type(s) used for the teacher. FSL: fully-supervised, WSL: weakly-supervised, SSL: self-supervised learning.
While early 3D CNN-based architectures can be referred to as foundation models [30], the term foundation model today commonly refers to Transformer-based [40] models that leverage large-scale pre-training. For vision, these models typically adopt the Vision Transformer [11] (ViT) architecture.
Unlike the convolutional architecture, which embeds strong spatial and temporal inductive biases, ViTs rely on learning such priors directly from data during pre-training [11]. The quality and scale of this pre-training directly influence their effectiveness in downstream tasks [50]. Pre-training methods vary in supervision type and scalability. Fully-supervised approaches rely on manually annotated labels, providing precise semantic guidance but limited scalability. Weakly-supervised methods, such as CLIP [33], use natural language or metadata as training signals. Though less curated, they offer rich semantic structure and broader concept coverage from web-scale data. Self-supervised learning (SSL) methods, including masked modeling [17, 38], learn from the data itself without any annotations, enabling large-scale pre-training and highly transferable representations. Unlike weak supervision, which relies on sparse and often noisy text labels [33], SSL provides denser and more unbiased training signals [6, 17].
First Video ViTs exploited supervised classification as their pre-training method. ViViT [1] adopted the ViT architecture to video by introducing spatiotemporal 3D cubes called tubelets instead of 2D patches used for images. While this demonstrated that attention-based models can handle video inputs, the method struggled to balance accuracy and efficiency. Subsequent models like TimeSformer [4], MViT [12], and VideoSwin [28] focused on improving efficiency.
VideoMAE [38] adopted masked autoencoding (MAE), a type of Masked Video Modeling (MVM), as an effective and efficient self-supervised pre-training strategy for plain video ViTs. Its tube masking, when applied to a large fraction of input patches, forces the model to infer spatiotemporal structure from limited visible content. This approach yielded strong results while maintaining architectural simplicity and has since inspired a series of ViT-based Video Foundation Models (ViFMs) that employ self-supervised pre-training [45, 44, 46].
While other architectures such as recurrent, hybrid, and state-space models are active research areas [23, 47, 31], at the moment, ViT-based ViFMs are arguably the dominant paradigm due to their strong performance, scalability, versatile pre-training strategies, and widespread optimization support (e.g., FlashAttention [9], optimized libraries, hardware acceleration) designed around the plain Transformer [11] architecture. These qualities, along with the growing availability of pre-trained Video ViTs, make them particularly promising for tasks like traffic anomaly detection (TAD), where generalization, robustness, and efficiency are critical.
To our knowledge, we are the first to research which type of video pre-training is most effective for the TAD task, and hypothesize that MVM, with its emphasis on learning patch-dense and temporally-aware representations, is particularly well-suited for this task.
3 Methodology
We fine-tune multiple ViT-based Video Foundation Models (ViFMs) for TAD and evaluate their performance against recent specialized TAD methods. We follow the encoder-only design and attach a single linear classification head to the output of the final encoder layer. This minimal design ensures that performance primarily reflects the effectiveness of the ViFM backbone in capturing patterns relevant for traffic anomaly detection. We investigate (i) whether a simple Video ViT model, pre-trained at scale, can achieve state-of-the-art performance on TAD, (ii) whether better general ViFMs are also better for TAD, and what type of pre-training is more effective (iii) and, finally, whether small-scale domain-adaptive pre-training (DAPT) is feasible and effective for adapting Video ViTs to the driving domain.
3.1 Task definition
We formulate Traffic Anomaly Detection (TAD) as a binary classification task, specifically focusing on frame-level, ego-centric anomaly classification, where each frame captured from a moving vehicle-mounted camera is assigned an anomaly label.
Let $X_{t}=\{I_{t-\tau+1},I_{t-\tau+2},\ldots,I_{t}\}$ denote a time-ordered sliding window of $\tau$ consecutive video frames captured from a vehicle-mounted camera up to time $t$. Each $I_{k}$ represents an RGB frame at time step $k$, from the egocentric viewpoint of the vehicle.
The task is to learn a function $f_{\theta}$ that maps an input window $X_{t}$ to a prediction $A_{t}$ at each timestep $t$:
$A_{t}=f_{\theta}(X_{t}),\quad t=T,T+1,\ldots$
where $A_{t}\in\{0,1\}$ is a binary label that indicates whether an anomaly is observed at time $t$.
In the general case, $\tau$ can be 1 and $f_{\theta}$ may also maintain an internal hidden state or operate in an autoregressive manner, explicitly conditioning on previously predicted outputs.
3.2 Evaluation Procedure
Prior work typically reports the Area Under the Receiver Operating Characteristic Curve (AUCROC) as the primary evaluation metric for TAD [48, 51, 24, 32, 25], and we adopt this metric when comparing to previous methods. However, handling data imbalance is especially important in TAD, so in our evaluations we use the Matthews Correlation Coefficient (MCC), which has been used in related work [42, 18, 37]. MCC takes into account all entries of the confusion matrix, including true negatives, and better reflects overall performance under class imbalance [7]. MCC at a given threshold is defined as:
$\text{MCC}=\frac{\text{TP}\cdot\text{TN}-\text{FP}\cdot\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}},$ (1)
where TP, TN, FP, and FN denote true positives, true negatives, false positives, and false negatives, respectively. Note that MCC ranges from $-1$ (inverse prediction) to $1$ (perfect prediction), with 00 indicating random performance, but we show it in the range -100, 100 to improve readability.
To assess discriminative ability independently of the decision threshold, we compute MCC across thresholds in the range $[0,1]$ and report the area under this curve, referred to as the Area Under the MCC Curve (AUCMCC). We also report MCC at a fixed threshold of 0.5 (MCC@0.5).
Beyond metric design, we implement a broader protocol focused on generalization and efficiency. We evaluate in-domain performance, out-of-domain performance, and computational cost.
3.3 Pre-trained Encoders
We select a range of recent ViFMs that represent various pre-training strategies, and apply them to the TAD task; see Tab. 1 for an overview of their pre-training strategies. When possible, we select variants pre-trained on Kinetics-400 [19] for consistency.
We include ViViT [1] as a baseline to represent fully-supervised pre-training. VideoMAE [38], MVD [45], and VideoMAE2 [44] are selected to evaluate progressively stronger variants of self-supervised pre-training from videos. We also assess InternVideo2 [46], which combines self-supervised learning from videos and weakly-supervised learning from multiple modalities, and is one of the leading models across numerous video benchmarks. Together, these ViT-based models cover a diverse range of pre-training strategies.
For completeness, we also include the fully-convolutional R(2+1)D model, pretrained in a fully-supervised manner, motivated by recent studies showing its competitive performance in the related task of collision anticipation [10].
3.4 Domain-Adaptive Pre-Training
To better align the Video ViT encoder with the driving domain, we adopt the Domain-Adaptive Pre-Training (DAPT) strategy, a simple method originally proposed in the field of natural language processing [16]. DAPT introduces an additional pre-training stage between generic pre-training and downstream fine-tuning, using unlabeled data from the target domain.
We apply VideoMAE-based [38] DAPT as follows:
  • Step 1: Generic pre-training. As before, we initialize the encoder with an off-the-shelf VideoMAE model pre-trained on large-scale generic video data, mostly unrelated to the driving domain.
  • Step 2: Domain-Adaptive Pre-training (DAPT). We continue pre-training the same model on a medium-sized dataset of unlabeled driving videos using the exact same VideoMAE reconstruction objective:
    $\mathcal{L}_{\text{DAPT}}=\left\|M\odot\left(x-f_{\theta}(x_{\text{masked}})\right)\right\|_{2}^{2}$
    where $x$ is the input video, $x_{\text{masked}}$ is the masked input, $f_{\theta}$ is the encoder-decoder VideoMAE model, $M$ is the binary mask, and $\odot$ is element-wise multiplication.
  • Step 3: Fine-tuning on TAD. As before, we fine-tune the encoder-only model on TAD datasets using the same configuration with a simple linear classification head.
The intermediate DAPT step (Step 2) specializes the model towards the driving domain without requiring any labels. It introduces no additional parameters, preserves model efficiency, and remains fully compatible with standard VideoMAE pipelines.
4 Experiments
DoTA AUCROC, % D2K AUCROC, %
[…]
Two-stage TAD methods
VidNeXt [10] 73.9 69.3 70.1 72.4 125 M 0.78 GB 27
ConvNeXt+VT [10] 73.1 61.2 66.8 67.3 125 M 0.77 GB 27
ResNet+NST [10] 74.0 70.1 71.2 72.3 24 M 0.19 GB 124
MOVAD [34] 82.2 77.6 77.0 75.2 153 M 1.10 GB 26
Multi-representation fusion TAD methods
TTHF [24] 84.7 71.7 140 M 0.80 GB 26
PromptTAD [32] 83.9 74.6 106 M 1.88 GB 18
ISCRTAD [25] 86.2 82.7 359 M 1.51 GB 33
Encoder-only models
R(2+1)D [39] 81.5 76.4 78.8 78.4 27 M 0.27 GB 104
DAPT-VideoMAE-S (ours) 86.4 81.7 85.6 84.3 22 M 0.16 GB 95
DAPT-VideoMAE-B (ours) 87.9 83.5 87.6 85.8 86 M 0.54 GB 94
DAPT-VideoMAE-L (ours) 88.4 84.2 88.5 86.6 304 M 1.80 GB 34
[…]
4.1 Experimental setup
Datasets.  We evaluate on DoTA [48] and DADA-2000 [13], two large-scale real-world driving anomaly datasets with temporal and frame-level annotations. DAPT uses Kinetics-700 [5], BDD100K [49], and CAP-DATA [14], described in detail in the Supplementary.
Model input.  All Video ViTs and R(2+1)D are trained on sliding windows of size 224×224×16 at 10 FPS (1.5s temporal context) by default. For InternVideo2, which uses tubelets of size 1, we use 224×224×8 at 5 FPS to match the same duration. MOVAD processes videos frame-by-frame at resolution 640×480.
Fine-tuning.  With all Video ViTs and R(2+1)D, we closely follow the VideoMAE fine-tuning recipe for HMDB51. We train for 50 epochs (5 warmup), with 50K randomly sampled examples per epoch and a batch size of 56. For VidNeXt variants, and MOVAD, we follow the original training settings.
Domain-adaptive pre-training (DAPT).  We apply the VideoMAE pre-training strategy [38], masking 75% of tokens, using MSE loss on masked tokens only. Training uses a batch size of 800 and 1M samples per epoch, with 12 epochs. We explore DAPT on three domains: (a) Kinetics-700, (b) BDD100K (normal driving), and (c) BDD100K + CAP-DATA (abnormal driving), with dataset mixing ratios detailed in the Supplementary.
4.2 Can an encoder-only model outperform specialized TAD methods?
To answer this question, we evaluate models along three critical axes: classification performance, generalization, and efficiency, as shown in Tab. 2. We select a range of recent top-performing methods proposed for the TAD task. Among encoder-only models, we apply the R(2+1)D [39] model and different sizes of VideoMAE pre-trained Video ViTs (with DAPT, see Sec. 4.4). The results show that these Video ViTs consistently strike a strong balance, demonstrating a good combination of predictive accuracy, generalization across domains, and computational efficiency. Notably, strongly pre-trained Video ViTs achieve the highest AUCROC scores across both DoTA and DADA-2000 datasets, both in-domain and in cross-dataset evaluation, while being highly efficient with a low memory footprint. In contrast, specialized TAD-specific models not only demonstrate lower classification performance but also incur substantially higher computational costs and latency. R(2+1)D and ResNet+NST, while being highly efficient, fall short in predictive quality. This confirms that we can outperform specialized, multi-component TAD methods with a simple encoder-only model by applying a Video ViT with strong pre-training.
4.3 What pre-training is better for TAD?
We investigate how general video recognition performance and pre-training strategies relate to downstream performance on TAD. We evaluate a range of publicly available Video ViT models of several sizes, using their Top-1 accuracy on Kinetics-400 [19] and Something-SomethingV2 [15] alongside AUCMCC on DoTA[48] and DADA-2000 [13].
Results are summarized in Tab. 3, from which we observe two key trends. First, we find that for TAD the MAE pre-training objective dominates: MAE-pre-trained models (VideoMAE) and their distilled variants (VideoMAE2, MVD) achieve the highest AUCMCC within each size tier, even when they do not have the highest classification accuracy on general benchmarks. Second, classification accuracy on general benchmarks is not representative of TAD performance: ViViT-B, despite matching VideoMAE on Kinetics-400, demonstrates significantly lower AUCMCC, and InternVideo2, state-of-the-art on general benchmarks, also underperforms on TAD. These findings suggest that the representations which are beneficial for general video classification may not align well with those needed for TAD. In particular, TAD appears to benefit more from dense representations that emphasize fine-grained temporal irregularities rather than the coarse semantic categories typically targeted by general video recognition models.
The overall top-ranking model on TAD is VideoMAE2, which incorporates dual masking, an additional pre-training step with distillation from a larger model, and $\sim$6 times larger-scale pre-training datasets, compared to other MVM pre-trained models. This confirms that both the scale of pre-training and the choice of objectives significantly impact the transferability of ViFMs to TAD.
Top-1 accuracy MCC@0.5 AUCMCC
Model Variant Type K400 SthSthV2 DoTA D2K DoTA D2K
VideoMAE1600 [38] Small SSL 79.0 66.8 55.5 49.5 52.1 48.1
MVDfromB [45] Small SSL 80.6 70.7 56.2 49.8 50.0 48.1
MVDfromL [45] Small SSL 81.0 70.9 56.5 51.1 50.2 49.1
VideoMAE2 [44] Small SSL+FSL 83.7 56.8 51.6 55.2 50.3
InternVideo2 [46] Small WSL+SSL 85.4 71.6 51.6 44.5 49.7 43.7
ViViT [1] Base FSL 79.9 30.7 27.6 28.9 26.7
VideoMAE800 [38] Base SSL 80.0 58.0 52.0 54.5 51.2
VideoMAE1600 [38] Base SSL 81.0 69.7 58.7 52.6 56.0 52.2
MVDfromB [45] Base SSL 82.7 72.5 57.8 51.6 56.0 50.9
MVDfromL [45] Base SSL 83.4 73.7 59.2 52.1 57.0 51.0
VideoMAE2 [44] Base SSL+FSL 86.6 75.0 58.4 54.8 56.5 53.4
InternVideo2 [46] Base WSL+SSL 88.4 73.5 52.2 44.2 50.0 43.1
VideoMAE1600 [38] Large SSL 85.2 74.3 61.6 56.9 59.7 55.36
MVDfromL [45] Large SSL 86.0 76.1 60.5 54.6 59.0 53.7
Table 3: Comparing Video ViT pre-trainings. In contrast to general video classification benchmarks (K400, SthSthV2), fully- and weakly-supervised pre-training are less effective for TAD benchmarks. Self-supervised pre-training performs best for TAD (DoTA, D2K). FSL: fully-supervised; WSL: weakly-supervised; SSL: self-supervised learning; K400: Kinetics-400 [19]; SthSthV2: Something-SomethingV2 [15]; D2K: DADA-2000 [13].
4.4 Domain-Adaptive Pre-Training (DAPT)
Larger ViFMs can be pre-trained on a larger scale and, as a result, exhibit better out-of-the-box generalization across domains, while smaller models have shown to benefit less from longer pre-training due to their limited capacity and faster saturation [50]. Therefore, we expect that domain adaptation can help better utilize the capacity of the smaller but at the same time more efficient and faster models. Given that MAE pre-training proves especially effective for TAD, and unlabeled driving data is available in abundance, we investigate whether small-scale self-supervised DAPT with MAE can be an effective and efficient way to scale the performance of smaller models. We initialize a ViT model with VideoMAE pre-trained weights and perform several epochs of additional pre-training with the VideoMAE objective on in-domain data. Compared to the original $\sim$192K training steps with batch size 2048, we use only 15K steps with batch size 800.
As shown in Fig. 3, DAPT via MAE brings clear improvements for small and base VideoMAE pre-trained models. As expected, given our small-scale DAPT protocol, the large model sees less improvement.
Figure 3: DAPT scaling across different model sizes. Smaller models benefit more. S: small, B: base, L: large variant of the Video ViT.
To disentangle the impact of domain relevance from that of additional pre-training, we conduct an ablation study on the data used for DAPT. Specifically, we adapt a model, pre-trained on a general human activity dataset, using three types of unlabeled video data: (1) the original, general pre-training domain used by VideoMAE, which is not related to the TAD task, (2) normal ego-centric driving, and (3) ego-centric driving mixed with anomalies. This setup allows us to evaluate whether the observed improvements stem from domain alignment or from simply continuing generic pre-training. As shown in Tab. 4, adaptation with domain-relevant data (both normal and abnormal driving) consistently improves generalization and data efficiency, while additional pre-training on the original, generic domain yields no notable gains. Interestingly, pre-training on normal driving videos is sufficient, and mixing in data with driving anomalies does not provide further improvements. We conclude that small-scale self-supervised DAPT is a simple and effective way to improve the performance and generalization of smaller Video ViTs for TAD, which does not necessarily requires rare anomaly data.
DoTA AUCMCC, D2K AUCMCC
[…]
w/o DAPT 52.1 43.8 48.1 46.6
Generic DAPT 51.6 -0.5 43.8 48.5 +0.4 46.2 -0.4
Driving DAPT 54.8 +2.7 46.8 +3.0 52.0 +3.9 49.7 +3.1
+ anomalies 54.9 +2.8 46.9 +3.1 51.9 +3.8 49.7 +3.1
[…]
Finally, in Fig. 4 we also include some qualitative examples which clearly demonstrate the positive effect of DAPT.
Figure 4: Qualitative examples for the effect of DAPT. Predicted anomaly‐scores of VideoMAE (top: S, bottom: B) w/ and w/o DAPT.
5 Discussion
In this work, we show that with stronger pre-training, an encoder-only Video Vision Transformer outperforms all prior Traffic Anomaly Detection models while also being significantly more efficient. However, it remains an open question whether the additional components introduced in earlier work become redundant as pre-training scales, as shown in related perception tasks [21], or whether they still provide complementary benefits.
We use Video Masked Autoencoders [38, 44] as a simple and effective form of masked video modeling (MVM). Approaches that predict in latent space, such as V-JEPA [3, 2], may offer further gains by avoiding the limitations of noisy pixel-level reconstruction.
6 Conclusion
Ego-centric Traffic Anomaly Detection (TAD) is a challenging task that requires modeling motion dynamics and agent interactions. While most recent methods for TAD rely on complex, multi-component architectures, we show that a simple encoder-only design using a plain Video Vision Transformer (ViT) with strong self-supervised pre-training is not only more efficient, but also more effective and generalizable. Building on this, Domain-Adaptive Pre-Training (DAPT) offers a label-free and data-efficient way to further boost performance, particularly for smaller models. These findings highlight the strength of learned inductive biases from large-scale pre-training as an alternative to manually crafted architectural complexity, a principle to which TAD is no exception.
Our experiments further demonstrate that Masked Video Modeling (MVM) is the most effective pre-training strategy for TAD, in contrast to standard video classification tasks. This suggests that different video tasks may benefit from pre-training objectives tailored to their downstream requirements. While TAD is a crucial task in autonomous driving (AD), other AD tasks may align more closely with conventional action recognition. This motivates further research into a universally effective video pre-training strategy, evaluated by its generalization across diverse AD tasks. We hope our findings provide a foundation for future work in this direction.
Acknowledgements
This work was funded by the Horizon Europe programme of the European Union, under grant agreement 101076754 (project AITHENA). Views and opinions expressed here are however those of the author(s) only and do not necessarily reflect those of the European Union or CINEA. Neither the European Union nor the granting authority can be held responsible for them. We also acknowledge the Dutch national e-infrastructure with the support of the SURF Cooperative, grant agreement no. EINF-10314, financed by the Dutch Research Council (NWO), for the availability of high-performance computing resources and support.